The pH of a 0.35 M solution of a weak base is 10.58. What is the Kb of the base?


1
Expert's answer
2020-12-01T09:03:32-0500

B(aq) + H2O(l) "\to" HB+ (aq) + OH- (aq)

I _____ 0.35M __________0____________ 0

C_______- x __________+ x __________ + x

E _____0.35M - x _________x____________x


pH = 10.58

[B] = 0.35M

kb = ?


pH = -log[H3O+]

[H3O+] = 1010.58 = 2.63×10-11M


kw = [H3O+][OH-]

[OH-] = "\frac{k_w}{[H_3O^+]}" , kW = 1.0×10-14 (Constant)


[OH-] = "\frac{1.0×10^-¹⁴}{2.63×10^-¹¹}" = 3.80×10-4 = x


[HB+] = 3.80×10-4 = x


Now we find kb

kb = "\frac{[HB^+][OH^-]}{[B]}"


kb = "\frac{(3.80×10^-⁴)(3.80×10^-⁴)}{0.35}"


kb = 4.12×10-7M